OpenStudy (anonymous):

one to one functions g and h are defined as follows. g={ (-8,6), (-4,5), (5,8), (9,3)} h (x)= 2x-3 and I have to find the following of: 1.) g^-1 (5) = ? 2.) h^-1 (x) =? (h o h^-1) (2)= ? (I don't understand the steps to finding these answers..)

jimthompson5910 (jim_thompson5910):

g^-1 (5) is asking what value of x will produce y = 5 when you plug it into g(x)?

jimthompson5910 (jim_thompson5910):

well that's x = -4 since we have the ordered pair (-4, 5) ie when x = -4, y = 5 in g(x) so, g^-1 (5) = -4

OpenStudy (anonymous):

ok i understand that one

jimthompson5910 (jim_thompson5910):

h^-1(x) is the inverse of h(x) so you need to find the inverse for part 2

jimthompson5910 (jim_thompson5910):

do you know how to do that?

OpenStudy (anonymous):

i see what your saying i don't know how to find the inverse of 2x-3 though..

jimthompson5910 (jim_thompson5910):

h(x) = 2x - 3 y = 2x - 3 ... replace h(x) with y x = 2y - 3 ... swap x and y now solve for y

OpenStudy (anonymous):

so would you add the 3 to x and divide both by 2?

OpenStudy (anonymous):

i got y=x/2+3/2

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so that's the inverse function h^-1(x) = x/2 + 3/2 you can optionally write it as h^-1(x) = (x + 3)/2

OpenStudy (anonymous):

im not sure what the last one is asking me to do

jimthompson5910 (jim_thompson5910):

(h o h^-1) (2) is the same as saying h( h^-1(2) )

OpenStudy (anonymous):

so am i plugging 2 into x

jimthompson5910 (jim_thompson5910):

you are plugging 2 into the inverse and getting some result...call this result z you then take this result and plug it into the function h(x) ie you are finding h(z)

OpenStudy (anonymous):

so im plugging 2.5 into h(x)?

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

so would 2 be the answer

jimthompson5910 (jim_thompson5910):

correct, it turns out that because h(x) is linear, this means h( h^-1(x) ) = x is true for all values of x

OpenStudy (anonymous):

ok thank you im starting to understand what these are asking!

jimthompson5910 (jim_thompson5910):

you're welcome